OpenStudy (anonymous):

A certain type of Swiss chocolate comes in a package that's the shape of a triangular prism. If the perimeter of one of the triangular bases is 16 centimeters, what is the surface area of the entire package? a. 164 cm2 b. 224cm2 c. 248 cm2 d. 272cm2

OpenStudy (anonymous):

do you have any questions about openstudy i can help with?

OpenStudy (anonymous):

Well, yeah.

OpenStudy (anonymous):

what are your questions

OpenStudy (anonymous):

A certain type of Swiss chocolate comes in a package that's the shape of a triangular prism. If the perimeter of one of the triangular bases is 16 centimeters, what is the surface area of the entire package? a. 164 cm2 b. 224cm2 c. 248 cm2 d. 272cm2

OpenStudy (anonymous):

Oops! I am so sorry. I was reading your question wrong!

OpenStudy (anonymous):

its ook

OpenStudy (anonymous):

Well. How do you ask a new question?

OpenStudy (anonymous):

in the math section? or another section

OpenStudy (anonymous):

In general

OpenStudy (anonymous):

Please someone I need help!

OpenStudy (anonymous):

what you need help with

OpenStudy (anonymous):

A certain type of Swiss chocolate comes in a package that's the shape of a triangular prism. If the perimeter of one of the triangular bases is 16 centimeters, what is the surface area of the entire package? a. 164 cm2 b. 224cm2 c. 248 cm2 d. 272cm2

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

Yep, thankyou so much.

terenzreignz (terenzreignz):

So many things to assume. One of them, hopefully, being the triangular base being an isosceles triangle.

terenzreignz (terenzreignz):

So, assuming that, the formula for the surface area is vague at worst, but it always involves getting the area of each of the faces of your figure, then adding them all up.

OpenStudy (anonymous):

I don't know how to do that.

OpenStudy (anonymous):

do you know the surface area formula

OpenStudy (anonymous):

I think the formula is A =1 over 2bh

OpenStudy (anonymous):

The formula for the SA is SA = 2B +(s1 +s2 +s3)h am I right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

What is it then?

OpenStudy (anonymous):

SA = wh + lw + lh + ls

OpenStudy (anonymous):

= width h = height l = length s = side

OpenStudy (anonymous):

is the width 6cm?

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

so SA = 6(4) + 14(6) +14(4) +14(?)

OpenStudy (anonymous):

What is the side?

OpenStudy (anonymous):

w=6,s=4,1=14

OpenStudy (anonymous):

SA = 6(4) + 14(6) +14(4) +14(5) is that correct?

OpenStudy (anonymous):

wait are you with me

OpenStudy (anonymous):

Yeah, I am solving for the formula.

OpenStudy (anonymous):

Y? am I doing it wrong?

OpenStudy (anonymous):

no i didnt give you a height yet

OpenStudy (anonymous):

Oh, lol, sorry.

OpenStudy (anonymous):

wait im doing it

OpenStudy (anonymous):

Doing what?

OpenStudy (anonymous):

the problem

OpenStudy (anonymous):

i work slow cause i get the right answer

OpenStudy (anonymous):

Area of two bases = 4 times 6 = 24 Area of bottom = 6 times 14= 84 Area of two sides = 2 times 5 times 14 = 140 Total area = 24+84+140=248

OpenStudy (anonymous):

you were partly right

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (anonymous):

np do you get it now

OpenStudy (anonymous):

yeah. Do you think you can answer one more question of mine?

OpenStudy (anonymous):

sure ask away

OpenStudy (anonymous):

What is the approximate length of the lateral surface for the cylinder shown here in net form? a. 31.4 in b. 62.8 in c. 78.5 in d. 376.8 in

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

the lateral surface of the cylinder=(area of the circle base) times (height of cylinder).

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

Yes. So, how do you find the area of a circle?

OpenStudy (anonymous):

the 10 is the diameter. The area of the circle base is pi*diameter of the circle

OpenStudy (anonymous):

31.4 is the area of the circle.

OpenStudy (anonymous):

So, 31.4 * 12 = 376.8

OpenStudy (anonymous):

correct good job your doing good

OpenStudy (anonymous):

anything else

OpenStudy (anonymous):

Yes, I think one more. (thank you soooo much, i really really appreciate it!!!)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

ask away

OpenStudy (anonymous):

What is the surface area this storage box? a. 15.625 ft ^2 b. 18.75 ft ^2 c. 37.5 ft ^2 d. 45 ft ^2

OpenStudy (anonymous):

So, 6(2.5) ^2 = 15

OpenStudy (anonymous):

wait no sry did it wrong hold on this is what happens when you dont take your time

OpenStudy (anonymous):

Lol, its okay, that is what I got at first, and I am stuck on it.

OpenStudy (anonymous):

each side is a square of length 2.5 and so area 2.52 there are 6 sides so you want 6×2.52

OpenStudy (anonymous):

Where did 2.52 come from, isn't it just 2.5?

OpenStudy (anonymous):

so it would be 2.5^2*6

OpenStudy (anonymous):

so its A anything else

OpenStudy (anonymous):

2.5 times 2.5=6.25 times 6

OpenStudy (anonymous):

I am just checking if this is right. I got 495 m^2. Calculate the surface area of this triangular prism. a. 255 m^2 b. 360 m^2 c. 468 m^2 d. 495 m^2

OpenStudy (anonymous):

=37.5

OpenStudy (anonymous):

how did you get that? and that is not in one of the a,b,c,d option.

OpenStudy (anonymous):

37.5ft^2 its C sry

OpenStudy (anonymous):

Ok, thank you.

OpenStudy (anonymous):

im doing the other one do you still need help or no

OpenStudy (anonymous):

and np

OpenStudy (anonymous):

Which other one?

OpenStudy (anonymous):

495m^2

OpenStudy (anonymous):

Oh, you were doing the box, I thought you were doing the triangle prism.

OpenStudy (anonymous):

thats the one im doing

OpenStudy (anonymous):

So the triangle prism is, 495m^2? and the box is 37.5 ft ^2?

OpenStudy (anonymous):

im not sure about 495m^2 but the rest are right since you didnt give me time to finish the other one

OpenStudy (anonymous):

Oh, sorry.

OpenStudy (anonymous):

well i got my own work to finish so yea

OpenStudy (anonymous):

just say yours is right

OpenStudy (anonymous):

sry bout that

OpenStudy (anonymous):

Its fine, you have been a lot of help! Thanks soo much again! Have a good rest of the day!

OpenStudy (anonymous):

you too bye